Shopping at Savers Mart, Lisa buys her children four shirts and three pairs of pants for $85.50. She returns that day and buys three shirts and five pairs of pants for $115.00. What is the price of each shirt and each pair of pants?
Answer by checkley77(12844)

4s+3p=85.50 multiply this equation by 3.
3s+5p=115.00 multiply this equation by -4 & add them.
12s+9p=256.50
-12s-20p=-460
-------------------
-11p=-203.5
p=-203.50/-11
p=$18.50 is the price of each pair of pants.
4s+3*18.50=85.50
4s+55.50=85.50
4s=85.50-55.50
4s=30
s=30/4
s=$7.50 is the price of each shirt.
Proof:
3*7.50+5*18.50=115.00
22.50+92.50=115.00
115.00=115.00